### v1

# Learning to Discover Generalized Facial Expressions

Abstract
We introduce Facial Expression Category Discovery (FECD), a novel task in the domain of open-world facial expression recognition (O-FER). While Generalized Category Discovery (GCD) has been explored in natural image datasets, applying it to facial expressions presents unique challenges. Specifically, we identify two key biases to better understand these challenges: Theoretical Bias—arising from the introduction of new categories in unlabeled training data, and Practical Bias—stemming from the imbalanced and fine-grained nature of facial expression data. To address these challenges, we propose FER-GCD, an adversarial approach that integrates both implicit and explicit debiasing components. In the implicit debiasing process, we devise F-discrepancy, a novel metric used to estimate the upper bound of Theoretical Bias, helping the model minimize this upper bound through adversarial training. The explicit debiasing process further optimizes the feature generator and classifier to reduce Practical Bias. Extensive experiments on GCD-based FER datasets demonstrate that our FER-GCD framework significantly improves accuracy on both old and new categories, achieving an average improvement of 9.8% over the baseline and outperforming state-of-the-art methods.
1 Introduction
Facial expression recognition (FER) is crucial in human-computer interaction, as expressions are a significant manifestation of human emotions [1, 2]. Traditional models are trained on seven basic facial expressions, i.e., happy, sad, surprise, anger, fear, disgust, and neutral. However, research indicates that humans can exhibit expressions that extend beyond these basic categories, including composite and complex expressions [3, 4], such as happy-surprise and perplexity. Consequently, models trained solely on these basic expressions perform poorly when encountering new types of expressions in real open-world FER (O-FER) scenarios. Additionally, manually annotating each expression in O-FER scenarios is prohibitively expensive and impractical. Consequently, reducing manual annotations and discovering generalized expression categories in O-FER have become important research focal points.
Recently, Open-Set FER [5, 6] has better solved the problem of models facing new expressions and is able to maintain high closed-set accuracy. However, current Open-Set FER methods still have limitations as they only aim at detecting new expressions without going for further new expression classification.
To further explore open-world scenarios and address previous limitations in O-FER, we introduced Generalized Category Discovery (GCD) [7] to FER scenarios for the first time, aiming to simultaneously recognize known expressions and discover different unknown expressions. Previous GCD approaches have made significant progress on general image datasets through research into contrastive learning techniques [7, 8, 9] and parametric classifier learning [10, 11].
However, directly applying existing GCD methods does not fully address the challenges outlined in this paper. These methods face inherent conflicts between recognizing previously learned categories and discovering new ones, resulting in biased learning. As training progresses, models tend to focus more on the new categories, causing a significant drop in accuracy for the old ones. Furthermore, the imbalanced distribution of expression data and the fine-grained differences between facial expressions exacerbate the bias, making it difficult for models to distinguish between subtle emotional variations. We summarize these challenges into two key points:
Figure 1: Left: GCD in O-FER aims to discover unkown (new), subtler facial expressions beyond the initial set of known (old) classes. Right: In GCD tasks, a major bias arises from new classes in unlabeled data interfering with model learning. In O-FER scenarios, this bias accumulates, sharply dropping old-class recognition accuracy in the mid-term. Our approach mitigates such biases and sustains a more stable improvement.
Challenge I: Theoretical Bias. In the semi-supervised GCD model, the introduction of unlabeled data with new emotion categories poses a challenge for the model’s learning, resulting in Theoretical Bias. During loss optimization, competition between old and new categories can lead to training bias, which is difficult to decouple in traditional GCD models with a single shared classification head, as shown in Fig. 9. Initially, the model improves in recognizing old categories due to its focus on labeled data. However, as training progresses, it increasingly prioritizes new categories, causing a significant drop in accuracy for the old categories, as shown in 1. This bias can be constrained through effective model design and optimization.
Challenge II: Practical Bias. In fine-grained and imbalanced O-FER scenarios, Practical Bias accumulates due to the characteristics of the data. Similarities between distinct expression categories and variations within the same category lead to misrecognition. The imbalanced data distribution favors majority expressions, resulting in lower accuracy for minority expressions. These accumulated biases hinder the model’s ability to capture distinguishing features, creating vague boundaries and difficulties in forming meaningful clusters.
Figure 2: Overview of FER-GCD framework. (i) In the implicit debiasing phase, we construct an adversarial process ($\mathcal{L}_{ad}$ ) between the Aux Head and the Main Head to help estimate the upper bound of bias and minimize it. (ii) In the explicit debiasing phase, we further enhance the performance of the feature generator $\psi$ and the Main Head through a combination of loss functions ($\mathcal{L}_{\text{rep}}$, $\mathcal{L}_{\text{cls}}$, $\mathcal{L}_{\text{bal}}$, $\mathcal{L}_{\text{cluster}}$). Clustering algorithm $\mathcal{L}_{\text{cluster}}$ will be introduced after the warm-up epochs.
In this paper, to address the challenges of Theoretical Bias and Practical Bias in O-FER, we propose a novel adversarial debiasing framework, named FER-GCD. FER-GCD integrates both implicit and explicit debiasing processes. The implicit debiasing process focuses on identifying and reducing the upper bound of Theoretical Bias, particularly caused by the introduction of new emotion categories. To achieve this, we define the F-discrepancy, which measures and minimizes the maximum bias in new categories. The explicit debiasing process addresses Practical Bias by enhancing the feature generator and classifier, leading to more discriminative feature boundaries that mitigate the effects of data imbalance and fine-grained variations. Together, these debiasing techniques significantly improve the robustness and accuracy of FER-GCD in tackling both types of bias in open-world FER scenarios.
Our contributions can be summarized as follows: (i) We introduce Facial Expression Category Discovery, aiming to understand human emotions in real and various scenarios. To the best of our knowledge, this is the first time to address this issue. (ii) We identify two main biases: Theoretical Bias, caused by new categories in unlabeled training data, and Practical Bias, arising from fine-grained and imbalanced facial expression scenarios. (iii) We propose FER-GCD, a novel adversarial framework integrating implicit and explicit debiasing techniques. (iv) Extensive experiments conducted on GCD-based FER datasets demonstrate approach’s superiority over other state-of-the-art GCD methods.
2 FER-GCD algorithm
In this section, we introduce the debiasing strategy of FER-GCD, which addresses both Theoretical Bias and Practical Bias through implicit and explicit debiasing. Specifically, in section 2.2, we define and constrain Theoretical Bias through mathematical bounds, followed by an adversarial training process for implicit debiasing. In section 2.3, we tackle Practical Bias through explicit debiasing, which enhances the feature generator and the main classification head to handle imbalanced and fine-grained facial expression data.
2.1 Preliminaries
Problem Setting. The GCD [7] problem involves maintaining the model’s ability to categorize known classes while also discovering new classes when given partially labeled data and a large amount of unlabeled data. In the context of FER datasets, we define the unlabeled dataset as $\mathcal{D}_{U}={(\mathbf{x}_{i}^{u},\mathbf{y}_{i}^{u})}\in\mathcal{X}\times\mathcal{Y}_{u}$, where $\mathcal{Y}_{u}$ is the label space for unlabeled data points. The goal of GCD in this context is to train a model that effectively categorizes the instances in $\mathcal{D}_{U}$ using information from a labeled dataset $\mathcal{D}_{L}={(\mathbf{x}_{i}^{l},\mathbf{y}_{i}^{l})}\in\mathcal{X}\times\mathcal{Y}_{l}$, where $\mathcal{Y}_{l}$ is the label space for labeled data points, and $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$. In this study, we assume the number of categories in the unlabeled space is known and is $\mathcal{Y}_{u}$, represented by $K_{u}=|\mathcal{Y}_{u}|$.
[…]
2.2 Implicit Debiasing for Theoretical Bias
Implicit debiasing targets Theoretical Bias, which arises from new categories in unlabeled data. We first estimate and constrain this bias, then minimize its upper bound through adversarial training.
Definition 1 (Metrics of bias)
With a sufficiently large amount of data, we define $\xi(\cdot,\cdot)$ as a measure of the difference between the model’s predictions and the ground truth, where the predictions go through a softmax layer, mapping to the probability space $\mathbb{P}$. A mapping function $f:\mathbb{P}\rightarrow\mathbb{E}$ then converts these probabilities into the Euclidean space $\mathbb{E}$. We can define the metric:
$\xi(\mathcal{H}_{h,\psi}(x),\mathcal{F}(x))=\sqrt{\sum_{i=1}^{n}\left|\right|\mathcal{H}_{h,\psi}(x_{i})-\mathcal{F}(x_{i})\left|\right|^{2}},$ (1)
The non-negativity, symmetry and metricity satisfied by $\xi(\cdot,\cdot)$ will help us to carry out the proofs that follow, please see the appendix 1 for the specific properties.
Since GCD models all use the semi-supervised training strategy, we define the prediction discrepancy of the models on labeled data $\mathcal{D}_{L}$ and unlabeled data $\mathcal{D}_{U}$ separately, we use the metric defined in defination1:
$\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})=\mathbb{E}_{x\in\mathcal{D}_{L}}\left[\xi\left(\mathcal{H}(x),\mathcal{F}(x)\right)\right],\quad\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})=\mathbb{E}_{x\in\mathcal{D}_{U}}\left[\xi\left(\mathcal{H}(x),\hat{\mathcal{F}}(x)\right)\right],$ (2)
Definition 2 (Metric promotion based on category space)
In GCD task, $\mathcal{D}_{U}$ contains both new and old categories that are inaccessible to us. The proportion of old and new classes can be assumed to follow a binomial distribution, where the probability of old class is denoted as $\theta$. Therefore, we can concretely define that the discrepancy on unlabeled data consists of the following convex combination:
$\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})=(1-\theta)\xi^{new}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})+\theta\xi^{old}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}}),$ (3)
Furthermore, since $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$, and based on empirical extrapolations, we can assume that the model will make more mistakes on unlabeled datasets (even if the categories are the same), mathematically represented as follows:
$\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})\leq\xi^{old}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}}),$ (4)
Definition 3 (F-discrepancy)
We define the upper bound on the discrepancy between the current model $\mathcal{H}$ and an arbitrary model hypothesis $\mathcal{H^{\prime}}$ on both labeled and unlabeled data. The F-discrepancy is:
$\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})=\sup_{\mathcal{H},\mathcal{H^{\prime}}\in\mathcal{R}}\left|\xi_{\mathcal{D}_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|,$ (5)
where $\alpha$ is a tuning parameter that adjusts the weight of the prediction bias for labeled data. The F-discrepancy metric is crucial for the subsequent lemma 1.
Lemma 1 (Upper Bound on Theoretical Bias from New Categories)
Since we cannot access the labels of $\mathcal{D}_{U}$ to determine whether the data belongs to new categories or old categories, we need to leverage $\mathcal{D}_{L}$ to help constrain the estimation of the upper bound of the new category bias in a rigorous manner. Eventually we get an upper bound on the bias of the new category:
$\xi^{new}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\frac{1}{1-\theta}\left[(\alpha-\theta)\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})+\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})+\lambda\right],$ (6)
Where $\lambda=\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H^{*}},\mathcal{F})+\xi_{\mathcal{D}_{U}}(\mathcal{H^{*}},\mathcal{F})$. Proof 2 is provided in the appendix.
Adversarial Optimization for Theoretical Bias: In Lemma 1, we establish an upper bound for the Theoretical Bias introduced by new categories in the unlabeled data. The goal of the implicit debiasing process is to minimize this upper bound. This serves to constrain the learning of new categories and helps avoid the pitfall of degrading the performance of previously learned categories due to the introduction of new data. This min-max process is computed as follows:
$\mathop{\min}_{\mathcal{H}_{h,\psi}}\mathop{\max}_{\mathcal{H^{\prime}}_{h_{a},\psi_{a}}}(\alpha-\theta)\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})+\left|\xi_{\mathcal{D}_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|,$ (7)
Where $(\alpha-\theta)$ is an adjustment parameter, $\psi$ represents the feature generator. Considering that in Definition 1, we use the mapping function $f$ to transform the softmax probability vector into Euclidean space, we now apply $f^{-1}:\mathbb{E}\rightarrow\mathbb{P}$ to revert back to the probability space. Thus, we use cross-entropy and construct the following adversarial loss:
Estimate the F-discrepancy, corresponding to the max process of Eq. (7).
$\mathcal{L}_{ad}=\frac{1}{n}\sum_{j=1}^{n}\ell_{CE}(h_{a}(\psi(x^{u}_{j}),\hat{\mathcal{P}}^{h}(x_{j}^{u}))-\frac{\alpha}{m}\sum_{i=1}^{m}\ell_{CE}(h_{a}(\psi(x^{l}_{i}),\mathcal{F}(x_{i}^{l})),$ (8)
Where $x^{u}$ and $x^{l}$ are labeled and unlabeled data, respectively, and $\hat{\mathcal{P}^{h}}$ is the pseudo-label assigned by main classification head $h$ to the unlabeled data, $\alpha$ coefficient is set to 2 according to the Tab 7. $\mathcal{L}_{ad}$ implies that approaching the bias requires $h$ and $h_{a}$ to be consistent on the labeled dataset $\mathcal{D}_{L}$, while being as inconsistent as possible on $\mathcal{D}_{U}$.
The implicit debiasing procedure has established the upper bound of bias and formulated an optimization target to minimize the maximum bias. The ’min’ component of Eq. (7) will be discussed in more detail during the explicit debiasing phase, corresponding to Eq. (12).
Figure 3: The ACC of single expression. Our approach effectively eliminates the bias caused by the introduction of new categories, and is able to better focus minority expressions (e.g., Ang., Neu.).
2.3 Explicitly Debiasing for Practical Bias
In the implicit debiasing phase, we establish a min-max optimization objective to address the Theoretical Bias of Challenge I. Additionally, for the Practical Bias of Challenge II, we aim to further enhance the performance of the feature generator and classifier to adapt to the O-FER scenario.
2.3.1 Optimization of the Feature Generator
Learning Subtle Facial Features aims to acquire discriminative features from feature generator $\psi$ that enable the classifier to effectively categorize all categories. In this stage, we utilize two types of contrastive learning. For any two random augmented versions $\hat{\mathbf{x}}_{i}$ and $\tilde{\mathbf{x}}_{i}$ of an image $\mathbf{x}_{i}$ within a training batch $\mathcal{B}$ that comprises both labeled and unlabeled samples. The self-supervised contrastive loss is defined as:
$\mathcal{L}_{\text{rep}}^{u}(\hat{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{i};\psi,\tau_{u})=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}-\log\frac{\exp(\cos(\psi(\hat{\mathbf{x}}_{i}),\psi(\tilde{\mathbf{x}}_{i}))/\tau_{u})}{\sum_{\mathbf{x}_{k}\in\mathcal{B}}\exp(\cos(\psi(x_{k}),\psi(\tilde{\mathbf{x}}_{i}))/\tau_{u})},$ (9)
where $\tau_{u}$ denotes a scaling parameter known as the temperature. Similar to Eq. (9), we construct supervised contrastive loss [12] $\mathcal{L}_{\text{rep}}^{s}(\hat{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{i},y;\psi,\tau_{c})$ on labeled data. These two losses are merged to shape the learning objective for the representation: $\mathcal{L}_{\text{rep}}=(1-\lambda)\mathcal{L}_{\text{rep}}^{u}+\lambda\mathcal{L}_{\text{rep}}^{s}$, with $\lambda$ serving as a tuning parameter.
Forming Discriminative Facial Expression Boundaries (Clustering). Clustering is essential for distinguishing expressions. To enhance feature discriminability, we integrate clustering with contrastive learning, which typically lacks global data structure awareness. Specifically, we connect the feature generator $\psi$ to a supervised clustering algorithm.
The clustering loss $\mathcal{L}_{\text{cluster}}$ comprises two components:
$\mathcal{L}_{\text{WB}}=\frac{\sum_{c\in\mathcal{C}^{l}}\sum_{x_{i}\in\mathcal{B}^{l}_{c}}\|\psi(x_{i})-\mu_{c}\|_{2}^{2}}{\sum_{c\in\mathcal{C}^{l}}n_{c}\|\mu_{c}-\mu_{g}\|_{2}^{2}+\epsilon},$ (10)
$\mathcal{L}_{\text{MM}}=\frac{1}{|\mathcal{C}^{l}|}\sum_{c\in\mathcal{C}^{l}}\left(\max_{x_{i}\in\mathcal{B}^{l}_{c}}\|\psi(x_{i})-\mu_{c}\|_{2}^{2}-\min_{x_{i}\in\mathcal{B}^{l}_{c}}\|\psi(x_{i})-\mu_{c}\|_{2}^{2}\right),$
The clustering loss is computed per mini-batch of labeled data $\mathcal{B}^{l}$, activated after $T_{\text{warmup}}$. Here, $\mu_{c}$ represents the class mean, $\mu_{g}$ is the global feature mean, and $n_{c}$ is the number of samples in class $c$. The set $\mathcal{C}^{l}$ includes all unique class labels in the batch, while $\mathcal{B}^{l}_{c}$ denotes the samples belonging to class $c$. The overall clustering loss is defined as $\mathcal{L}_{\text{cluster}}=\mathcal{L}_{\text{WB}}+\beta\cdot\mathcal{L}_{\text{MM}}$. The term $\mathcal{L}_{\text{WB}}$ aims to minimize intra-class distances and maximize inter-class distances, while $\mathcal{L}_{\text{MM}}$, as a regularization term, promotes compact clustering by reducing the gap between maximum and minimum distances.
2.3.2 Fine-tuning of Main Classification Head
Parametric classifier learning has become a effective paradigm. We use it to augment main head $h$. We follow the SimGCD [10] approach to assign labels to instance inputs. Specifically, the number of categories $K=\mathcal{M}+\mathcal{N}$ is given. A set of parametric prototypes for each category $\mathcal{T}=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{K}\}$ is randomly initialized at the beginning. During training, the soft label $\hat{\mathbf{p}}_{i}^{k}$ for each augmented view $\mathbf{x}_{i}$ is calculated using a softmax function on the cosine similarity between the hidden feature and the prototypes:
$\hat{\mathbf{p}}_{i}^{k}=\frac{\exp\left(\frac{1}{\tau_{s}}(h(\psi(\mathbf{x}_{i}))/{\|h(\psi(\mathbf{x}_{i}))\|_{2}})^{\top}({\mathbf{t}_{k}}/{\|\mathbf{t}_{k}\|_{2}})\right)}{\sum_{j}\exp{\left(\frac{1}{\tau_{s}}(h(\psi(\mathbf{x}_{i}))/\|h(\psi(\mathbf{x}_{i}))\|_{2})^{\top}(\mathbf{t}_{j}/\|\mathbf{t}_{j}\|_{2})\right)}},$ (11)
Similarly, we can obtain the soft label $\tilde{\mathbf{p}}_{i}$ of the view $\tilde{\mathbf{x}}_{i}$. The supervised and unsupervised losses of the classifier are formulated by:
$\mathcal{L}_{\text{cls}}^{s}$ $=\frac{1}{|\mathcal{B}^{l}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}^{l}}\ell_{CE}(\mathbf{y}({x_{i}}),\hat{\mathbf{p}}_{i}),\,\mathcal{L}_{\text{cls}}^{u}=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}\ell_{CE}(\tilde{\mathbf{p}}_{i},\hat{\mathbf{p}}_{i})-\epsilon H(\overline{\mathbf{p}}),$ (12)
where $\mathcal{B}^{l}$ is the mini-batch of labeled training data, $\mathbf{y}({x_{i}})$ is the ground truth label for the labeled data point $\mathbf{x}_{i}$, $\ell_{CE}$ is the cross-entropy loss, and $H(\overline{\mathbf{p}})=-\sum\overline{\mathbf{p}}\log\overline{\mathbf{p}}$ regularizes the mean prediction $\overline{\mathbf{p}}=\frac{1}{2|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}(\hat{\mathbf{p}}_{i}+\tilde{\mathbf{p}}_{i})$ in a mini-batch. Then the objective of parametric classifier learning is $\mathcal{L}_{\text{cls}}=(1-\lambda)\mathcal{L}_{\text{cls}}^{u}+\lambda\mathcal{L}_{\text{cls}}^{s}$.
Elimination of imbalance bias. In FER scenarios, sample imbalance often biases models toward majority classes, reducing accuracy for minority classes. This issue is further complicated when the minority class distribution is unknown. To address this, we introduce a dynamic reconciliation process that adaptively adjusts training, improving model generalization. We achieve this by utilizing the following loss function:
$\mathcal{L}_{\text{bal}}(\mathbf{x},y;\psi,h)=-\sum_{x_{i}\in\mathcal{B}}\sum_{c\in\mathcal{C}}\delta(y_{i},c)\left[(1-e_{i})+\frac{e_{i}}{a_{c}+\epsilon}\right]\log\left(\frac{\exp(h(\psi(\mathbf{x}_{i}))_{c})}{\sum_{k\in\mathcal{C}}\exp(h(\psi(\mathbf{x}_{i}))_{k})}\right),$ (13)
Where $a_{c}$ is the adaptive weight for class $c$, calculated as the ratio of $\varsigma_{c}$ (the number of correct predictions for class $c$) to $\eta_{c}$ (the total number of predictions made for class $c$), i.e., $a_{c}=\frac{\varsigma_{c}}{\eta_{c}}$. The time-varying weight $e_{i}$ decreases as training progresses, given by $e_{i}=0.1\cdot\left(1-\frac{t}{T}\right)$, where $t$ is the current training round and $T$ is the total rounds. This method dynamically tracks poorly predicted categories, eliminating the need for prior knowledge about the number of classes.
Overall loss function: Through the aforementioned process, we have defined the maximum bias and enhanced the model’s robustness in the presence of the maximum bias to eliminate its impact. The overall loss function is as follows:
$\mathcal{L}_{total}=\mathcal{L}_{rep}+\mathcal{L}_{cls}-\lambda_{a}\mathcal{L}_{ad}+\lambda_{b}\mathcal{L}_{bal}+\lambda_{c}\mathcal{L}_{cluster}$ (14)
Where $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{c}$ are three balancing parameters, we set them to 0.2, 0.3, and 0.2, respectively, based on the hyperparameter analysis in appendix B.3. $\mathcal{L}_{cluster}$ is only introduced after the warmup epoch $T_{warmup}$, when the model has preliminary clustering capabilities.
Table 1: Category discovery accuracy (ACC) on RAF-DB containing only known expressions. We test Ada-CM under GCD setting. $\triangle$ represents the improvement accuracy.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
k-means [13] 27.8 28.5 27 26.2 23.9 26.4 26.2 25.7 27.5
GCD [7][CVPR’22] 31.2 33.3 28.5 33.2 36.8 32.6 56.8 63.5 36.1
GPC [8][ICCV’23] 47.2 56.7 38.4 45.5 47.3 43.2 59.4 67.1 38.4
InfoSieve [14][NeurIPS’23] 52.8 72.2 40.1 50.1 62.2 48.3 58.2 64.6 39.6
SimGCD [10][ICCV’23] 43.3 47.9 40.2 49.6 56.9 48.5 57.1 62.1 41.7
Ours 53.3 71.3 41.0 52.4 66.3 50.2 69.1 77.8 41.9
Ada-CM [15][CVPR’22] 11.7 18.2 7.3 7.5 10.4 3.9 12.4 9.7 14.8
Ada-CM+ our debias framework 17.4 29.2 9.4 12.3 16.4 7.4 14.5 11.4 18.6
$\triangle$ +5.7 +11.0 +2.1 +4.8 +6.0 +3.5 +2.1 +1.7 +3.8
Table 2: Category discovery accuracy (ACC) on FerPlus.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
k-means [13] 20.8 22.1 20.2 22.5 23.0 22.3 25.1 25.6 24.0
GCD [7][CVPR’22] 40.8 62.3 27.4 40.4 52.2 32.2 37.2 50.4 29.7
GPC [8][ICCV’23] 45.2 63.7 34.4 45.7 64.3 40.0 52.2 58.5 46.8
InfoSieve [14][NeurIPS’23] 48.2 74.1 37.7 47.5 72.5 44.1 64.1 69.5 53.7
SimGCD [10][ICCV’23] 45.8 65.4 38.3 46.4 70.2 42.2 63.7 67.2 55.0
Ours 51.8 84.9 38.8 48.6 81.7 43.5 68.6 74.3 56.4
Table 3: Category discovery accuracy (ACC) on large-scale dataset AffectNet.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
k-means [13] 14.4 14.6 14.3 15.0 20.3 13.2 14.2 15.2 13.8
GCD[7][CVPR’22] 35.0 47.5 28.1 32.2 42.0 28.9 28.0 38.2 24.6
GPC [8][ICCV’23] 36.3 46.5 33.4 38.7 50.5 32.1 33.6 44.1 30.0
InfoSieve [14][NeurIPS’23] 47.1 54.2 35.7 49.8 68.4 37.2 46.2 57.3 31.4
SimGCD [10][ICCV’23] 45.6 53.4 35.1 40.7 55.1 34.5 38.1 48.1 29.7
Ours 57.7 62.9 40.9 54.2 74.8 39.4 55.9 69.4 30.8
3 Experiments
3.1 Experimental Setup
Dataset. We conduct experiments on three popular FER datasets, which have been partitioned in the format suitable for GCD. RAF-DB [16]contains eleven compound expressions and seven basic expressions. For GCD, we treat four out of the seven basic expressions as known categories, with three different selection methods. Notably, in the stage of discovering compound expressions, we use the basic expressions as known categories. FERPlus [17] is extended from FER2013 [18]. We used eight categories and selected four as known categories to discover the remaining new categories. AffectNet [19] is a large-scale FER dataset. We used eight categories and selected four as known categories to discover the remaining new categories. A summary of dataset statistics is shown in the appendix B.1. Please refer to the appendix for a statistical summary of the dataset and the reasons for the specific category partitioning.
Evaluation Metric. Following GCD [7] guidelines, we evaluate model performance using Category discovery accuracy (ACC). The Hungarian algorithm [20] is used to optimally assign emerged clusters to their ground truth labels. The ACC formula is given by $\text{ACC}=\frac{1}{Z}\sum_{i=1}^{Z}\mathbb{I}(y^{t}_{i}=q(\hat{y}_{i}))$, where $Z=|\mathcal{D}_{U}|$, representing the total number of samples in the unlabeled dataset, and $q$ is the optimal permutation that best matches the predicted cluster assignments to the ground truth labels. We report the accuracies of all the classes (“All”), old classes (“Old”) and unseen classes (“New”).
Implementation details. For a fair comparison, we ran our method and the comparison methods five times on each dataset, and ultimately reported the set of results where the "All" accuracy ranked third. We employ a ViT-B/16 backbone network [21] pre-trained with DINO [22]. Training was performed using an initial learning rate of 0.1, which was decayed with a cosine annealed schedule [23] . The max training epoch $T$ is set to 200 and batch size of 128 for training. We follow SimGCD [10] to set the balancing factor $\lambda$ to 0.35, and the temperature values $\tau_{c}$ and $\tau_{u}$ to 0.1 and 0.07, respectively. We initially set $\tau_{t}$ to 0.07 and $\tau_{s}$ to 0.1 for the classification objective. Then, a cosine schedule is employed to gradually reduce $\tau_{t}$ to 0.04 over the first 30 epochs. $\lambda_{a}$, $\lambda_{b}$, and $\lambda_{c}$ are assigned values of 0.2, 0.3, and 0.2, respectively. $T_{warmup}$ is set to 50 epochs. The coefficients $\alpha$ and $\beta$ are set to 0.2 and 0.1, respectively. The All experiments are conducted using an NVIDIA GeForce RTX 3090 GPU.
Figure 4: Comparison with open-set FER [5] in the open-set scenario.
Table 4: Ablation studies of different components of our method on the RAF-DB basic expression classes, showing both implicit and explicit debiasing effects.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang.
Method All Old New All Old New
SimGCD (baseline) 43.3 47.9 40.2 49.6 56.9 48.5
w/ $\mathcal{L}_{\text{ad}}$ (implicit debias only) 50.4+7.1 63.2+15.3 40.5+0.3 50.5+0.9 60.5+3.6 48.6+0.1
w/o $\mathcal{L}_{\text{ad}}$ (explicit debias only) 49.8+6.5 61.1+13.2 40.7+0.5 51.5+1.9 63.2+6.3 50.3+1.8
w/o $\mathcal{L}_{\text{cluster}}$ (implicit + $\mathcal{L}_{\text{bal}}$) 52.3+9.0 65.3+17.4 40.4+0.2 50.7+1.1 63.4+6.5 47.5-1.0
w/o $\mathcal{L}_{\text{bal}}$ (implicit + $\mathcal{L}_{\text{cluster}}$) 51.4+8.1 64.2+16.3 40.7+0.5 50.7+1.1 61.5+4.6 48.9+0.4
Ours (all components) 53.3+10.0 71.3+23.4 41.0+0.8 52.4+2.8 66.3+9.4 50.2+1.7
Figure 5: Attention visualization of different heads (numbered as h1 to h3) on RAF-DB. The top 10% attended patches are shown in red. Our method pays more attention to the cheeks, eyes and mouth corners details. Our model learns more discriminative features than other methods.
3.2 Comparison with the State-of-the-Art methods
Comparision with FER method. Given that the previous FER methods were not tailored for the GCD task, we opted for Ada-CAM [15], a semi-supervised approach, as a comparative baseline. We subjected Ada-CAM to our GCD evaluation metrics, and the results, as shown in Tab. 1, underscore the superiority of our method in the category discovery task. Furthermore, in light of the recent emergence of the open-set scenario for FER, introduced by [5] et al., we adapted the evaluation metrics for FER-GCD to align with the Open-set paradigm. Specifically, we utilized AUROC (higher is better) and FPR@TPR95 (lower is better) as our evaluation metrics. As shown in Fig. 4, FER-GCD achieved higher AUROC metrics, representing the greater capability for new category detection. In addition, the low FPR@TPR95 also reflects the confidence level of FER-GCD in making decisions, further reflecting the effectiveness of the debiasing framework.
Comparision with GCD methods. The effectiveness of FER-GCD debiasing is most directly reflected in its ability to maintain high accuracy for old facial expression categories while also achieving excellent results in discovering new facial expression categories. In Tab. 1, we present a comparison with the SOTA GCD method on the basic classes of the RAF-DB dataset. In terms of old category accuracy, we achieved an average improvement of 16.2% compared to SimGCD [10]. And the ability to discover new expressions has increased steadily. In Tab. 6, we report the comparison on the challenging RAF-DB-Compound dataset that discovers composite expressions based on basic expressions, we achieved an improvement of 20.8% accuracy for the old categories and 4.9% for the new categories, compared to SimGCD. Moreover, Tab. 2 shows the results on the dataset FerPlus [17]. Compared to SimGCD, we achieved significant improvements, whereas InfoSeive [14] is able to achieve an even higher accuracy for new categories when Surprise, Fear, Disgust, and Anger were known categories. Furthermore, Tab. 3 shows performance comparison on the large-scale dataset AffectNet [19]. On this challenging dataset, we both achieve the best new category, old category accuracy. In summary, our approach effectively eliminates the bias towards new categories while maintaining high accuracy for old categories, and steadily improves the recognition accuracy for new categories. An important basis for the outstanding performance of our method in identifying old categories and discovering new categories is that it forms more discriminative feature boundaries, as illustrated in Fig. 8.
3.3 Ablation Study
In this section, we analyze the importance of the components we introduce mainly on the RAF-DB basic classes.
Effect of estimating the maximum bias (implicit debias). We first examine the role of estimating maximum bias, employing an auxiliary classification head for adversarial estimation to ensure robust performance of the feature extractor under maximum bias. Tab.  4 presents the ablation results. We strive to minimize interference from new category introduction on old categories while improving accuracy for new categories.
Enhancing the effect of clustering (explicit debias). In FER-GCD, we introduce the clustering algorithm after the warmup epoch $T_{\text{warmup}}$, assuming that the model has preliminary clustering capabilities by then. This allows us to further constrain the feature space through our clustering algorithm. Tab.  4 presents the ablation results.
Focus on minority categories (explicit debias). A smoother dynamic focus mechanism allows us to focus well on a small number of categories and further improves overall accuracy. Tab.  4 demonstrates the validity of this technique.
4 Visualisation
As shown in Fig. 5, we adopt the method from GCD [7] to conduct a visual analysis of the attention maps generated by the DINO-ViT model. Specifically, we visualize the attention Heads 1 through 3, which provides valuable insights into the crucial aspects of the model’s recognition process, particularly how it captures the subtle nuances in facial expressions. Besides, the t-SNE visualization results shown in Fig. 8 demonstrate a more discrete feature representation of FER-GCD, compared to the more blurred feature boundaries of GCD and SimGCD.
5 Conclusion
In this paper, we introduce Generalized Category Discovery (GCD) for facial expression recognition (FER), with the aim of reducing manual annotation costs and improving the model’s ability to detect new compound expressions, thereby enhancing future sentiment analysis. During training, interference between old and new classes can bias the model, especially in fine-grained and imbalanced FER scenarios. To address this, we first theoretically defined the bias and proposed a new framework called FER-GCD, which debiases through implicit and explicit steps. Finally, our method is able to maintain good performance in recognizing old expressions while enhancing the ability to discover new expressions. Our work aims to advance facial expression recognition for open-world tasks.
Appendix A Theory
A.1 Notation
[…]
A.2 Metric of Bias
Definition 1. With a sufficiently large amount of data, we define $\xi(\cdot,\cdot)$ as a measure of the difference between the model’s predictions and the ground truth. where the model’s predictions go through a softmax layer. Since there is a sufficient amount of data, we use the following metric:
$\xi(\mathcal{H}_{h,\psi}(x),\mathcal{F}(x))=\sqrt{\sum_{i=1}^{n}\left|\right|\mathcal{H}_{h,\psi}(x_{i})-\mathcal{F}(x_{i})\left|\right|^{2}},$ (15)
Next we need to prove that it satisfies the metric properties.
Proof 1
To simplify the calculations:
$\mathcal{H}_{h,\psi}(x)=\mathcal{H}(x)\And\mathcal{H^{*}}_{h^{*},\psi^{*}}(x)=\mathcal{H^{*}}(x),$ (16)
$\|\mathcal{H}(x)-\mathcal{F}(x)\|^{2}=\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\|^{2}+\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\|^{2}+2(\mathcal{H}(x)-\mathcal{H^{*}}(x))\cdot(\mathcal{H^{*}}(x)-\mathcal{F}(x)),$ (17)
$\sum_{x=1}^{\left|D\right|}\left|\right|\mathcal{H}(x)-\mathcal{F}(x)\left|\right|^{2}$ (18)
$=\sum_{x=1}^{\left|D\right|}\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\|^{2}+2\sum_{x=1}^{\left|D\right|}(\mathcal{H}(x)-\mathcal{H^{*}}(x))\cdot(\mathcal{H^{*}}(x)-\mathcal{F}(x))+\sum_{x=1}^{\left|D\right|}\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\|^{2},$
Since the cross terms may be negative, we need to safely estimate the upper bound, according to Cauchy-Buniakowsky-Schwarz Inequality:
$\left[\sum_{x=1}^{\left|D\right|}(\mathcal{H}(x)-\mathcal{H^{*}}(x))\cdot(\mathcal{H^{*}}(x)-\mathcal{F}(x))\right]^{2}\leq\left[\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\right\|^{2}\right]\cdot\left[\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\right\|^{2}\right],$ (19)
$\left|\sum_{x=1}^{\left|D\right|}(\mathcal{H}(x)-\mathcal{H^{*}}(x))\cdot(\mathcal{H^{*}}(x)-\mathcal{F}(x))\right|\leq\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\right\|^{2}}\cdot\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\right\|^{2}},$ (20)
This allows us to perform a safe deflation to estimate the upper bound:
$\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H}(x)-\mathcal{F}(x)\right\|^{2}$ (21)
$=\sum_{x=1}^{\left|D\right|}\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\|^{2}+2\sum_{x=1}^{\left|D\right|}(\mathcal{H}(x)-\mathcal{H^{*}}(x))\cdot(\mathcal{H^{*}}(x)-\mathcal{F}(x))+\sum_{x=1}^{\left|D\right|}\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\|^{2}$
$\leq\sum_{x=1}^{\left|D\right|}\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\|^{2}+2\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\right\|^{2}}\cdot\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\right\|^{2}}+\sum_{x=1}^{\left|D\right|}\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\|^{2},$
Finally we get the inequality:
$\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H}(x)-\mathcal{F}(x)\right\|^{2}}$ (22)
$\leq\sqrt{\sum_{x=1}^{\left|D\right|}\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\|^{2}+2\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\right\|^{2}}\cdot\sqrt{\sum_{x=1}^{\left|D\right|}\left\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\right\|^{2}}+\sum_{x=1}^{\left|D\right|}\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\|^{2}}$
$=\sqrt{\sum_{x=1}^{\left|D\right|}\|\mathcal{H}(x)-\mathcal{H^{*}}(x)\|^{2}}+\sqrt{\sum_{x=1}^{\left|D\right|}\|\mathcal{H^{*}}(x)-\mathcal{F}(x)\|^{2}},$
So we have:
$\xi(\mathcal{H}(x),\mathcal{F}(x))\leq\xi(\mathcal{H}(x),\mathcal{H^{*}}(x))+\xi(\mathcal{H^{*}}(x),\mathcal{F}(x)),$ (23)
And it’s clear that:
$\xi(\mathcal{H}(x),\mathcal{F}(x))\geq 0\And\xi(\mathcal{H}(x),\mathcal{F}(x))=\xi(\mathcal{F}(x),\mathcal{H}(x)),$ (24)
A.3 Bounding the Bias
Definition 3 (F-discrepancy) For the hypothesis space $\mathcal{R}$, we define the upper bound on the discrepancy between the current model $\mathcal{H}$ and an arbitrary model hypothesis $\mathcal{H^{\prime}}$ on both labeled and unlabeled data. The F-discrepancy is:
$\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})=\sup_{\mathcal{H},\mathcal{H^{\prime}}\in\mathcal{R}}\left|\xi_{\mathcal{D}_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|,$ (25)
where $\alpha$ is a tuning parameter, and in fact it is possible to change the weight of the prediction bias for labeled data by changing $\alpha$. Next we need to use F-dicrepancy in the proof of lemma 1.
Lemma 1 (Bounding New Category Bias). Let $\mathcal{R}$ be a hypothesis space under the Fer-GCD training process. $\mathcal{H}$, $\mathcal{H^{\prime}}$, and $\mathcal{H^{*}}$ represent the current model, any arbitrary model hypothesis, and the model that minimizes the joint error, respectively. In order not to access the labels of $\mathcal{D}_{U}$, we need to constrain them with labels on the $\mathcal{D}_{L}$ and theoretically define the upper bound of bias for new categories:
[…]
Proof 2
According to Definition 1, we have:
[…]
So we have:
[…]
$D_{U}$ contains both new and old categories:
[…]
According to Definition 2, we have:
$\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})\leq\xi^{old}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}}),$ (30)
Considering that we want to constrain the new categories of bias:
$\xi_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})=(1-\theta)\xi^{new}_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})+\theta\xi^{old}_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\xi_{D_{U}}(\mathcal{H},\mathcal{H^{*}})+\xi_{D_{U}}(\mathcal{H^{*}},\hat{\mathcal{F}}),$ (31)
Adding and subtracting terms:
$(1-\theta)\xi^{new}_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})+\theta\xi^{old}_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\xi_{D_{U}}(\mathcal{H},\mathcal{H^{*}})+\xi_{D_{U}}(\mathcal{H^{*}},\hat{\mathcal{F}})-\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{*}})+\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{*}}),$ (32)
Rearrange the equation:
$(1-\theta)\xi^{new}_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\xi_{D_{U}}(\mathcal{H^{*}},\hat{\mathcal{F}})-\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{*}})+\xi_{D_{U}}(\mathcal{H},\mathcal{H^{*}})+\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{*}})-\theta\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F}),$ (33)
Based on $\xi(\cdot,\cdot)$ metric properties, further scaling is applied:
$(1-\theta)\xi^{new}_{D_{U}}(\mathcal{H},\hat{\mathcal{F}})$ (34)
$\leq\xi_{D_{U}}(\mathcal{H},\mathcal{H^{*}})-\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{*}})+\alpha\cdot\xi_{D_{L}}(\mathcal{H^{*}},\mathcal{F})+\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})-\theta\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})+\xi_{D_{U}}(\mathcal{H^{*}},\hat{\mathcal{F}})$
$\leq\left|\xi_{D_{U}}(\mathcal{H},\mathcal{H^{*}})-\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{*}})\right|+(\alpha-\theta)\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})+\lambda$
$\leq\underbrace{\sup_{\mathcal{H},\mathcal{H^{\prime}}\in\mathcal{R}}\left|\xi_{D_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|}_{\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})}+(\alpha-\theta)\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})+\lambda,$
where the constant $\lambda=\alpha\cdot\xi_{D_{L}}(\mathcal{H^{*}},\mathcal{F})+\xi_{D_{U}}(\mathcal{H^{*}},\hat{\mathcal{F}})$.
Appendix B Experiments
B.1 Dataset Details
We adopt the class division from [24] utilizing 50% of the images from these labeled classes as labeled instances in $D_{L}$. The remaining images from these classes are considered as the unlabeled data $D_{U}$. RAF-DB [16] contains eleven compound expressions and seven basic expressions.
Table 5: Specific category divisions.
Old Classes New Classes
RAF-DB-Basic $|\mathcal{Y}_{l}|=4$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=3$
RAF-DB-Compound $|\mathcal{Y}_{l}|=7$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=11$
FerPlus $|\mathcal{Y}_{l}|=4$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=4$
AffectNet $|\mathcal{Y}_{l}|=4$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=4$
For GCD, we treat four out of the seven basic expressions as known categories, with three different selection methods. Notably, in the stage of discovering compound expressions, we use the basic expressions as known categories.
The RAF-DB-Basic dataset is focused on basic expressions, while the RAF-DB-Compound dataset includes more complex expressions. We selected a subset of the basic expressions as old classes for the purpose of our experiments. FERPlus [17] is an extension of the FER2013 dataset [18]. We selected a subset of the categories as old classes to discover the remaining new categories in this dataset. AffectNet [19] is a large-scale facial expression recognition dataset. Similar to the selection strategy for FERPlus, we selected four categories as old classes in AffectNet to discover the remaining expressions.
Table 6: Category discovery accuracy (ACC) on RAF-DB-Compound. We utilize the seven basic expressions training to discover new compound expressions.
Method All Old New
k-means [13] 15.7 16.6 13.9
GCD [7][CVPR’22] 22.0 26.4 20.1
GPC [8][ICCV’23] 25.2 28.1 22.3
InfoSieve [14][NeurIPS’23] 40.1 50.5 24.2
SimGCD [10][ICCV’23] 32.6 39.0 24.8
Ours 49.6 59.8 29.7
The basis for partitioning. We used the following partitioning means on the three datasets (RAF-DB-Basic, FerPlus, AffecNet): (Sur.+Fea.+Dis.+Hap.) as known categories represent a more balanced partitioning. Relatively, (Sur.+Fea.+Dis.+Ang.) represents more new categories and (Sur.+Hap.+Sad.+Neu.) represents more old categories. Please note that this is just one of the many approaches we have adopted for partitioning, and there are indeed numerous other ways to do so. In real-world deployments, partitioning can actually be done randomly.
B.2 Different Backbones
We explored the effect of three backbone networks ResNet-18, ResNet-50 and ViT-B-16 (the main tabular results presented in the paper) on RAF-DB. In fact, Vaze et al. [7] have explored the effect of different backbone networks on other datasets such as CIFAR10 and CIFAR100. We found that ResNet-18 and ResNet-50 are terrible at discovering new expressions in FER scenarios. On the clustering task, they perform poorly. As shown in Fig. 1, the accuracy decreases in the early stage, so the statistics of the final results are not very meaningful.
Figure 6: It can be old that using ResNet-18 and ResNet-50 as the backbone network will give poor results and the accuracy will show a decreasing trend, so the statistics on the final accuracy are not very meaningful.
B.3 The Influence of the Hyperparameters
Our method has three main hyperparameters, which are the weight of Loss Functions. We show the results of our model in Fig. 7 on RAF-DB by tuning $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{c}$ with different values. In general our model has good stability for different hyperparameters. We note that if too much weight is applied to the adversarial bias loss, the model performance decreases. Since we only use the adversarial loss to estimate the upper bound of bias, paying excessive attention to this loss can make the model difficult to optimize by other components. In addition $\mathcal{L}_{bal}$ should preferably not be weighted more than the contrastive learning loss $\mathcal{L}_{rep}$, which can lead to limitations in the model’s ability to learn feature representations. Furthermore, $\mathcal{L}_{cluster}$ aims to further improve the expression-discriminability of the feature boundaries and we find that the weight that most improves model recognition is 0.2. Here, we have only roughly obtained an optimal combination of loss weights without conducting a precise analysis. Our main focus was on exploring the effectiveness of the components. The hyperparameters $\alpha$ and $\beta$ are analyzed as shown in Tab. 7 and Fig. 7, taking values of 0.2 and 0.1, respectively.
Table 7: Category discovery accuracy (ACC) on RAF-DB. The effect of the hyperparameter $\alpha$
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
Ours($\alpha$=1) 52.6 70.4 40.5 51.5 62.6 49.8 68.9 78.0 40.8
Ours($\alpha$=2, version in the paper) 53.3 71.3 41.0 52.4 66.3 50.2 69.1 77.8 41.9
Ours($\alpha$=3) 53.0 72.0 40.4 51.8 66.5 49.2 68.7 77.2 41.7
Ours($\alpha$=4) 52.4 73.4 38.2 50.6 67.5 47.9 67.9 76.3 40.4
Figure 7: Hyperparametric analysis. $\lambda_{a}$, $\lambda_{b}$, and $\lambda_{c}$ represent the weights of the three losses, and $\beta$ represents the coefficients of the regular terms in the $\mathcal{L}_{\text{cluster}}$.
B.4 Comparison on general image datasets
In addition to the validation on the FER dataset, the results of our method on the natural image scenes CIFAR-10 and Herbarium 19 are shown in Tab 8. It is worth noting that Herbarium 19 is an unbalanced scene, whereas our method is more advantageous.
Table 8: Results on two general image recognition tasks, namely Herbarium 19 and CIFAR10.
Herbarium 19 CIFAR10
Methods All Old New All Old New
$k$-means 13.0 12.2 13.4 83.6 85.7 82.5
RS+ 27.9 55.8 12.8 46.8 19.2 60.5
UNO+ 28.3 53.7 14.7 68.6 98.3 53.8
ORCA 20.9 30.9 15.5 81.8 86.2 79.6
GCD 35.4 51.0 27.0 91.5 97.9 88.2
SimGCD 44.0 58.0 36.4 97.1 95.1 98.1
Ours 45.2 58.7 37.9 97.0 95.3 97.8
B.5 Sensitivity of the number of categories
In this study, we use a parametric classifier approach assuming that the categories are known [10, 25]. In this section, we analyze the effect of different number of categories on ACC. Tab 9 shows the results, and since FER is a highly imbalanced scenario, increasing the number of output categories of the output header appears as a rise in the old categories. In future work, we will further explore methods that do not have a number of categories a priori.
Table 9: Effect of different number of clusters in our FER-GCD on RAF-DB.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
category number=5 55.9 72.3 44.9 52.0 66.5 49.4 68.6 78.0 39.4
right category number 7 53.3 71.3 41.0 52.4 66.3 50.2 69.1 77.8 41.9
category number=10 50.5 72.8 35.4 50.4 63.1 48.5 68.9 78.5 39.1
category number=20 44.3 77.3 22.0 24.0 67.8 16.9 64.7 75.4 31.7
B.6 Improvement of Clustering Effect
In the GCD task, whether a category is correctly identified lies in whether a discriminative cluster is formed. Generating discriminative feature spaces becomes particularly difficult in fine-grained and imbalanced O-FER scenarios. Specifically we compare three approaches: (a) Running k-means [13] directly after processing features in DINO [22]. (b) GCD [7]. (c) SimGCD [10]. As can be old in Fig. 8, our method eliminates the bias well and further improves the discriminability of the feature clustering boundaries through the clustering algorithm.
Figure 8: T-SNE visualization of representations on RAF-DB. Our method produces more discriminative feature boundaries that help the model perform better expression recognition.
Appendix C GCD Model with Shared Classification Head
We mentioned in the main text that the introduction of new categories in unlabelled data leads to bias accumulating on a single shared categorical head. Whereas past GCD models have often used a single shared classification head, they are unable to decouple this bias.
Figure 9: An example of the previous GCD model, other method components may exist, but most are in the form of a single shared classification head. This configuration makes it challenging to prevent the accumulation of bias, ultimately leading to incorrect facial expression recognition by the model.
Appendix D Pseudo Code
We summarize the pipeline of FER-GCD in Algorithm 1.
Algorithm 1 The proposed FER-GCD framework
Input: Train set $\mathcal{D}=\{\mathcal{D}_{L}\cup\mathcal{D}_{U}\}$, feature generator $\psi$, main head $h$, auxiliary head $h_{a}$, train epoch $T$, warm-up epochs $w$.
Output: Trained model parameter $\mathcal{S}$.
Initialize: Load DINO  [22] pre-trained parameters for the backbone.
  for $\text{epoch}=0,\cdots,T-1$ do
     Calculate the predictions of the main head with Eq. 12 and Eq. 13;
     Compute the supervised loss $\mathcal{L}_{\text{rep}}^{s}(\mathcal{D}_{L})$ and self-supervised loss $\mathcal{L}_{\text{rep}}^{u}(\mathcal{D})$ with Eq. 9;
     Feed $\mathcal{D}$ to $h(\psi(\cdot))$ and $h_{a}(\text{GRL}(\psi(\cdot)))$GRL: Gradient Reversal Layer}
     Calculate the Adversarial Loss $\mathcal{L}_{ad}$ with Eq. 8;
     if $epoch\geq w$ then
        Feed $\mathcal{D}$ to $\psi(\cdot)$ and then obtain features;
        Compute the Cluster Loss $L_{cluster}(\psi(\mathcal{D})$;
     end if
  end for
After training stage, we would typically perform the GCD [7] metric evaluation directly in the program. The test strategy is summarized in Algorithm 2.
Algorithm 2 The test stage strategy of FER-GCD
Input: The test set $\mathcal{D}^{t}=\left\{(\boldsymbol{x}_{j},{y}_{j})\right\}\in\mathcal{X}\times\mathcal{Y}$, trained FER-GCD model $f_{s}$.
Output: Classification accuracy $\mathrm{ACC}$.
  for $\boldsymbol{x}_{i}\in\mathcal{D}^{t}$ do
     Obtain the feature of $\boldsymbol{x}_{j}$ via $f_{s}(\boldsymbol{x}_{j})$
  end for
  Calculate the optimal assignment between clusters and categories by Hungarian algorithm  [20];
  Compute the test accuracy $\mathrm{ACC}$ based on the optimal assignment.
Appendix E Related Work
Facial expression recognition(FER) conveys rich information and is a key focus of current AI research [26, 27, 28, 29, 30, 1]. There have also been many studies in recent years aimed at improving the performance of FER models [31, 32, 15, 33]. For instance, Yang et al. [31] introduce a de-expression learning method. Wang and his team propose an effective Self-Cure Network (SCN). Besides, semi-supervised learning [15] and challenging imbalance problems [33] are also important research directions in FER. Recently, in order to further explore open-world settings, Zhang et al. [5] introduce open-set setting to FER scenario. In addition, Liu et al. proposed video-baset open-set setting in FER [6]. However the open-set task aims to detect new classes that do not belong to the previous known category without further categorization of the new class, which remains a limitation.
Semi-supervised learning (SSL) utilize labeled and unlabeled data in the training stage, which is an effective method when labeled training data is scarce. Methods such as pseudo-labeling and consistency regularization, which are well-established in this field, have been shown to enhance performance [34, 35]. Moreover, FixMatch [36] and FlexMatch [37] improved the reliability of pseudo labelling by introducing a confidence-based thresholding technique. Furthermore, introducing contrastive learning methods based on SimCLR [38] and MoCo [39] in SSL can further enhance the representation learning ability.
Adversarial training [40] can help semi-supervised learning framework to generate some fake samples [41, 42, 43] or adversarial samples [44, 45]. Some works also use adversarial training to estimate worst-case scenarios [46, 47]. In our study, we theoretically bound the bias by the introduction of new categories and use adversarial training to estimate this maximum bias.
Category discovery aims to discover new categories from unlabeled data. During training, labeled and unlabeled training data provide important visual conceptual information. The assumption of Novel category discovery (NCD) [48, 49, 50, 51] is that there is no overlap in categories between the labeled and unlabeled datasets. However, this assumption is unrealistic in the broader open-world [7, 9, 52] setting, where the unlabeled dataset not only encompasses categories previously learned by the model from labeled data but also novel categories. Prior studies have devised efficient techniques for category discovery, utilizing parametric classifiers [10], enhancing representation learning [9, 53], or employing prompting learning with larger models [52]. In addition, some researches aim to focus on more realistic or difficult GCD problems, such as incremental learning-GCD [54], long-tail recognition GCD [55], cross-domain GCD [56], etc. However, most existing GCD methods utilize a single shared classification head, thus failing to decouple the accumulation of biases.
Appendix F Limitations and Future Work
Currently, FER-GCD deployments are primarily confined within the same domain, overlooking the challenge of domain shift or cross-domain scenarios. However, real-world environments are far more intricate, encompassing a wider range of complex and nuanced expressions. Moreover, we did not conduct a thorough parameter tuning, and the auxiliary head may not necessarily be the optimal construction method, so the model may not achieve the best possible performance. In the future, we intend to redirect our focus towards more open-ended scenarios, particularly the problem of cross-domain FER-GCD.